How fast should you drive if you want to travel 400 miles in 6 hours? (The distance formula is D=rt?

if possible could you explain this problem step by step how to do it so i will understand how to do it?

Found 2 solutions by Alan3354, RAY100:
Answer by Alan3354(69443)

How fast should you drive if you want to travel 400 miles in 6 hours? (The distance formula is D=rt?
--------------
d = rt
distance = rate * time
400 = rate * 6
rate = 400 miles/6 hours = 200/3 miles/hours (or mph)
rate = 200/3 mph = 66.67 mph
----------------
Notice that the units, miles and hours, are also involved. miles/hour is common, which implies that we divide the # of miles by the # of hours.

Answer by RAY100(1637)

Velocity = Distance / time = 400 miles / 6 hrs = 66.67 miles per hr
.